# Binary population synthesis of the Galactic canonical pulsar population

### v1

@@ 4 Future Development
    Some NSs are formed through accretion-induced collapses, and these NSs are not accounted for in this work. This is mainly because it takes much longer than 200 Myrs for a NS to form through this channel, and they are naturally left out from this study as we only choose to include binary systems formed within the most recent 200 Myrs. Additionally, neutron stars formed through accretion-induced collapse may directly form MSPs (e.g., Hurley et al., 2010; Tauris et al., 2013; Ruiter et al., 2019), which are not the focus of this work.
@@
    We currently lack a physical model for the birth spin periods and magnetic fields for pulsars as a function of the progenitor properties (cf. Mandel & Müller, 2020, for neutron star masses and kicks). This work, and any follow-up in the foreseeable future will only be able to use some empirically assumed distribution of these quantities.
@@
    We have neglected the population of magnetars with long spin periods and high magnetic fields (e.g., Popov et al., 2010; Gullón et al., 2015; Makarenko et al., 2021). As with canonical pulsars, most magnetars are isolated, whilst their massive star progenitors were presumably born in binaries. Binary interactions such as stellar mergers may be responsible for producing highly magnetic neutron stars (Schneider et al., 2019).
@@
    A full scope of pulsar population synthesis can enable the study of NS contribution to the diffuse $\gamma$-ray background (Fermi-LAT Collaboration, 2017)
@@
    The inclusion of MSPs in the population synthesis models would provide further insight into the possible MSP origin of the Galactic Centre GeV Excess (Hooper & Mohlabeng, 2016; Ploeg et al., 2017; Eckner et al., 2018).
@@
    A comparison of pulsar populations between the Galactic field and globular clusters with results from dynamical simulations, such as those from NBODY and CMC (e.g., Ye et al., 2019, 2024).
@@
    After adding in the formation of MSPs and NSs formed through accretion-induced collapse, the model will be able to perform on the Hubble time scale. This will enable the modelling of merger events for gravitational wave detections, such as black hole-NS binaries (Chattopadhyay et al., 2021), double NSs (Chattopadhyay et al., 2020) and NS-white dwarf binaries.
@@
    Include magnetar physics in future versions of COMPAS and synthesize magnetar populations.
@@
    Investigate more sophisticated model comparison methods to fully explore the parameter space and identify a best-fit model that produces the population of pulsars observed to date. Several approaches have been taken in the literature, including hierarchical Bayesian parameter estimation (e.g., Ploeg et al., 2020) and machine learning approaches (e.g., Ronchi et al., 2021; Graber et al., 2023). At present, such a large parameter space exploration is computationally infeasible using COMPAS. In a follow-up study, we will optimise, parameterise and automate our COMPAS simulations and post-processing pipelines so that the use of more sophisticated model exploration and comparison techniques becomes computationally feasible.
In addition to varying the birth distributions of the pulsar properties (birth magnetic fields and spin periods), in future work, we aim to study how different binary properties, including initial orbital properties, different mass transfer theories, neutron star kick velocities and common envelope physics can impact the pulsar population, and how this population can be used to constrain the underlying physics of massive binary evolution.
@@ 5 Conclusions
We performed a population synthesis of the canonical Galactic pulsar population using the rapid binary population synthesis code COMPAS (Riley et al., 2022) on both radio and $\gamma$-ray canonical pulsars. We generated 20 different models for pulsar birth properties for COMPAS runs incorporating a magnetic braking model with a braking index of 3 for pulsar spindown. We discover that the PN1-B1 model, where the birth period of pulsar follows a normal distribution with $\mu=75$ ms and $\sigma=25$ ms, the birth magnetic field of pulsars follows a uniform distribution between $10^{11}$ and $10^{13}$ G, and magnetic decay time scale $\tau_{d}=500$ Myr best describes the observed pulsar population, based on the facts that:
@@
    PN1-B1 model produces the number of radio and $\gamma$-ray canonical pulsars that, given the star formation rate of the model, best fits the catalog.
@@
    The distribution of physical properties of radio and $\gamma$-ray pulsars produced from this model matches the catalog the best.
We then compare in detail the PN1-B1 model results to the catalog by examining the P-$\dot{P}$ plots, the Galactocentric coordinates, and the heuristic flux between the simulation and the catalog. We also discussed the impact of the selection effect on the population synthesis results, demonstrating that the slightest change in pulsar physics and observation sensitivity can bring significant differences to the final simulation.
Using the results from PN1-B1 model, we also attempted to explore the possible emission mechanism from low $\dot{E}$ pulsars ($\dot{E}$ $<10^{33}$ erg s-1, and failed to make a distinction between a beamed emission mechanism or an isotropic emission mechanism.
@@
The authors appreciate coding assistance from R. Wilcox and J. Riley, and helpful discussions from T. Paglione, J. Tan and J. Hurley.
@@
This research was conducted by the Australian Research Council Centre of Excellence for Gravitational Wave Discovery (OzGrav), through project numbers CE170100004 and CE230100016. S. Stevenson is a recipient of an ARC Discovery Early Career Research Award (DE220100241). D. Chattopadhyay is supported by the UK’s Science and Technology Facilities Council grant ST/V005618/1. This work used the OzSTAR high-performance computer at Swinburne University of Technology. OzSTAR is funded by Swinburne University of Technology and the National Collaborative Research Infrastructure Strategy (NCRIS).
@@
COMPAS is open source and hosted on Github. Due to the size of COMPAS output files, the authors can share configuration files used to produce the pop-synth results upon request.
Post-processing scripts will be hosted on Github as well. Before being made available on Github, they will be available upon request provided that this work is properly cited.

### v2

@@ 4 Future Development
    Some NSs are formed through accretion-induced collapses, and these NSs are not accounted for in this work. We exclude them as neutron stars formed through accretion-induced collapse may directly form MSPs (Hurley:2010MNRAS; Tauris:2013A&A; Ruiter:2019MNRAS, e.g.,), which are not the focus of this work. Our models also do not include neutron stars formed from stellar merger products. These populations, whilst not expected to produce the majority of neutron stars, may go some way to explaining the low pulsar birth rates found in this study.
@@
    Our model currently assumes a binary fraction of 100% for massive stars. Whilst this is appropriate for massive O-type stars Sana:2012Sci; Sana:2014ApJS, B-type stars likely have somewhat lower binary fractions around 70% MoeDiStefano::2017ApJS; Zapartas:2017A&A. Future work should incorporate a mass dependent binary fraction and the contribution of single stars. This will be particularly important hen considering millisecond pulsars and accretion induced collapses, as binary interactions play a more important role in the evolution of these systems.
@@
    Our model of an exponentially decaying magnetic field (Section 2.3) is meant to provide a simple starting point for exploring the possibility of pulsar magnetic fields decaying on a broad range of timescales, as well as allowing for comparison to previous work Kiel:2008MNRAS; Chattopadhyay2020; Dirson2022. Recent work has introduced more sophisticated models for magnetic field decay motivated by detailed theoretical work (Vigano:2021CoPhC; Sarin:2023ApJL; Graber:2024ApJ, e.g.,). We leave the implementation of these models, and an exploration of their impact to future work.
  5. 5.
    We currently lack a physical model for the birth spin periods and magnetic fields for pulsars as a function of the progenitor properties (Mandel:2020qwb, cf.). This work, and any follow-up in the foreseeable future will only be able to use some empirically assumed distribution of these quantities.
  6. 6.
    We have neglected the population of magnetars with long spin periods and high magnetic fields (Ferrario:2006MNRAS; Ferrario:2008MNRAS; Popov:2010MNRAS; Gullon:2015MNRAS; Makarenko:2021MNRAS, e.g.,). As with canonical pulsars, most magnetars are isolated, whilst their massive star progenitors were presumably born in binaries. Binary interactions such as stellar mergers may be responsible for producing highly magnetic neutron stars Ferrario:2009MNRAS; Schneider:2019Nature; Frost:2024Sci.
  7. 7.
    Our model assumes that the angle between the magnetic and spin axes of pulsars is constant for all pulsars and does not evolve with time. This is a simplification, and we will explore the implications of a varying angle in future work (Johnston2020, e.g.,).
  8. 8.
    Updated treatment of both radio and $\gamma$-ray selection effects, including the dependence of pulsar radio luminosity and beam geometry on pulsar spin properties (Posselt:2023MNRAS, cf.) will improve the fidelity of our models.
@@
    A full scope of pulsar population synthesis can enable the study of NS contribution to the diffuse $\gamma$-ray background 2017arXiv170500009F.
@@
    The inclusion of MSPs in the population synthesis models would provide further insight into the possible MSP origin of excess of GeV $\gamma$-ray emission from the Galactic Centre hooperm2016; ploeg2017; eckner2018; Gautam:2022NatAs.
@@
    A comparison of pulsar populations between the Galactic field and globular clusters with results from dynamical simulations, such as those from NBODY simulations (Ye:2019ApJ; Ye:2024ApJ, e.g.,).
@@
    After adding in the formation of MSPs and NSs formed through accretion-induced collapse, the model will be able to perform on the Hubble time scale. This will enable the modelling of merger events for gravitational wave detections, such as black hole-NS binaries Chattopadhyay2021, double NSs Chattopadhyay2020 and NS-white dwarf binaries.
@@
    Including stellar mergers and the physics of magnetar formation and evolution in future versions of COMPAS will allow us to synthesise magnetar populations and compare them to observations. This will open up avenues for comparing COMPAS simulations to observations of fast radio bursts.
@@
    Investigate more sophisticated model comparison methods to fully explore the parameter space and identify a best-fit model that produces the population of pulsars observed to date. Several approaches have been taken in the literature, including hierarchical Bayesian parameter estimation (Ploeg:2020JCAP, e.g.,) and machine learning approaches (Ronchi:2021ApJ; Graber:2024ApJ; PardoAraujo:2025A&A, e.g.,). At present, such a large parameter space exploration is computationally infeasible using COMPAS. In a follow-up study, we will optimise, parameterise and automate our COMPAS simulations and post-processing pipelines so that the use of more sophisticated model exploration and comparison techniques becomes computationally feasible.
In future work, we aim to study how different binary properties, including initial orbital properties, different mass transfer theories, neutron star kick velocities and common envelope physics can impact the pulsar population, and how this population can be used to constrain the underlying physics of massive binary evolution.
@@ 5 Conclusions
We performed a population synthesis of the canonical Galactic pulsar population using the rapid binary population synthesis code COMPAS COMPAS on both radio and $\gamma$-ray canonical pulsars. We generated 40 different models for pulsar birth properties for COMPAS runs incorporating a magnetic braking model with a braking index of 3 for pulsar spindown. We discover that the PN1-B1-D4 model, where the birth period of pulsar follows a normal distribution with $\mu=75$ ms and $\sigma=25$ ms, the birth magnetic field of pulsars follows a uniform distribution between $10^{11}$ and $10^{13}$ G, and magnetic decay time scale $\tau_{d}=1$ Myr best describes the observed pulsar population, based on the facts that:
@@
    PN1-B1-D4 model produces the number of radio and $\gamma$-ray canonical pulsars that, given the star formation rate of the model, best fits the catalogue.
@@
    The distribution of physical properties of radio and $\gamma$-ray pulsars produced from this model matches the catalogue the best.
We then compare in detail the PN1-B1-D4 model results to the catalogue by examining the P-$\dot{P}$ plots, the Galactocentric coordinates, and the heuristic flux between the simulation and the catalogue. We also discussed the impact of the selection effect on the population synthesis results, demonstrating that the slightest change in pulsar physics and observation sensitivity can bring significant differences to the final simulation.
Using the results from PN1-B1-D4 model, we also attempted to explore the possible emission mechanism from low $\dot{E}$ pulsars ($\dot{E}$ $10^{33}$ erg s-1. Our results seem to favour the isotropic emission mechanism over the beamed emission mechanism.
@@
The authors appreciate coding assistance from R. Wilcox and J. Riley, and helpful discussions from J. Hurley.
@@
This research was conducted by the Australian Research Council (ARC) Centre of Excellence for Gravitational Wave Discovery (OzGrav), through project numbers CE170100004 and CE230100016. S. Stevenson is a recipient of an ARC Discovery Early Career Research Award (DE220100241). D. Chattopadhyay is supported by the UK’s Science and Technology Facilities Council grant ST/V005618/1. This work used the OzSTAR and Ngarrgu Tindebeek (NT) high-performance computers at Swinburne University of Technology. OzSTAR and NT are funded by Swinburne University of Technology and the National Collaborative Research Infrastructure Strategy (NCRIS), and the Victorian Higher Education State Investment Fund (VHESIF).
@@
COMPAS is open source and hosted on Github (https://github.com/TeamCOMPAS/COMPAS). Due to the size of COMPAS output files, the authors can share configuration files used to produce the pop-synth results upon request.
Our post-processing scripts are available on Github at https://github.com/yuzhesong/compas_pulsar_paper.git.